ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

5. Differential Geometry and Global Analysis

1991 MS Classification: 53 Kashani, S. M. B. ,Tarbiat Modarres University, Iran Cohomogeneity One Revolution Hypersurfaces of the Euclidean Space In this paper we study some immersed (noncompact) hypersurfaces of the Euclidean space which are cohomogeneity one Riemannian manifolds , i. e. acted on by a compact connected Liesubgroup of intrinsic isometries with principal orbits of codimension one. We give necessary and sufficient conditions in order that such a hypersurfaces be a “revolution hypersurface”.